I am having trouble solving this equation its so simple but i cant remember please help.
5x-3y=4
y-x=-2
solving by substitution method please show me how to solve I am not getting the answer right.

Answer by checkley71(8403)

X-Y=-2 OR X=Y-2 THEN SUBSTITUTING (Y-2) FOR X IN THE FIRST EQUATION WE GET
5(Y-2)-3Y=4 OR 5Y-10-3Y=4 OR 2Y=4+10 OR 2Y=14 OR Y=14/2 OR Y=7
THEN X-7=-2 OR X=-2+7 OR X=5
PROOF 5(7-2)-3*7=4 OR 5*5-21=4 OR 25-21=4 OR 4=4